# Rational Tuning of LLM Cascades via Probabilistic Modeling

### v3

@@ 4.4 Rational Tuning of Confidence Thresholds
Table 7: Contrasts the normalized area under the cost-error curve (AUC) on the test set, showing that our continuous optimization-based algorithm for selecting confidence thresholds consistently outperforms grid search, by an average of 1.0%. The mean percentage change (%$\%\Delta$) is statistically significant at the $p<0.05$ level on five out of six benchmarks, as measured by a Wilcoxon rank-sum test paired by cascade.
@@
(b) The runtime of our method always grows linearly in the desired resolution $h$ at which the cost-error curve is sampled along the cost axis, independent of the cascade length $k$. However, the runtime of grid search scales as $h^{k-1}$ in $h$ and $k$ (log-log plot).
@@
Figure 6: Shows runtime scaling for computing the full cost-error curve, comparing our continuous-optimization based algorithm (“continuous”, blue) to grid search (“grid search”, gray). Our method scales much more favorably as the cascade length grows, and as the cost-error curve is sampled more densely along the cost axis. The shading shows ±$\pm 1\sigma$ of the observed data points.
In this section, we examine the performance and runtime scaling of our continuous optimization-based algorithm (11) for selecting optimal confidence thresholds. We consider all 26 possible cascades of length $k\geq 2$ composed of Meta’s Llama models (1B, 3B, 8B, 70B, and 405B) and evaluate against a grid search baseline on six benchmarks (MMLU, MedMCQA, XSum, TriviaQA, GSM8K, TruthfulQA) spanning general-purpose knowledge and reasoning, domain-specific QA, text summarization, open-ended QA, mathematical reasoning, and the ability to avoid hallucinations on adversarial questions.
Performance metrics: we evaluate the area under the cost-error curve (AUC) on the test set. Specifically, computing the AUC means plotting the test error ($y$ axis) against the expected cost ($x$ axis) and evaluating the integral of this curve. We normalize the cost values to lie between 0 and 1, resulting in AUC scores between 0 and 1. Lower AUC is better, as it indicates fewer mistakes at the same cost. In addition, we measure how the runtime for finding optimal confidence thresholds scales with the length of the cascade and the desired resolution of the cost-error curve on the $x$ axis, i.e., how densely we sample the optimal thresholds. We have not overly optimized our code and mainly aim to contrast asymptotic scaling behavior.
Grid search baseline: this baseline selects optimal confidence thresholds by searching over adaptive grids computed from the model-specific quantiles of calibrated confidence. Specifically, in each dimension the grid ranges from $\phi_{\text{min}}$ to $\phi_{\text{max}}$ in increments of $2.5\%$ probability mass. After scoring all candidate threshold combinations, we use the Skyline operator (implemented in the Python package paretoset [^5] Open-source implementation available at github.com/tommyod/paretoset. Accessed January 13, 2025.) to filter the candidate threshold vectors down to the Pareto-optimal set (Börzsönyi et al.,, 2001). A candidate threshold vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{k-1})$ is Pareto-optimal if its performance metrics $(\mathbb{P}_{\boldsymbol{\theta}}(\text{Correct}),\mathbb{E}_{\boldsymbol{\theta}}[\text{Cost}])$ are not dominated by any other candidate threshold vector $\boldsymbol{\theta^{\prime}}$ in the sense that $\mathbb{P}_{\boldsymbol{\theta^{\prime}}}(\text{Correct})>\mathbb{P}_{\boldsymbol{\theta}}(\text{Correct})$ and $\mathbb{E}_{\boldsymbol{\theta^{\prime}}}[\text{Cost}]<\mathbb{E}_{\boldsymbol{\theta}}[\text{Cost}]$.
Figure 5 shows that our continuous optimization based-method for selecting confidence thresholds results in lower AUC on the test set, compared to grid search. Each point on the plot shows the mean change in AUC for all cascades of a given length $k$, averaged across all benchmarks. As cascade length $k$ grows, our method outperforms by a larger margin. For example, the mean reduction in AUC for $k\geq 2$ is 1.9%; $k\geq 3$, it is 2.2%, and for $k=5$ it is 2.6%. We computed statistical significance of these percentage differences using a Wilcoxon rank-sum test paired by cascade. Interpreting these results, it is not surprising that grid search struggles as $k$ increases since searching for effective threshold combinations by trial and error suffers from the curse of dimensionality.
Table 7 presents these same results broken down by benchmark rather than cascade length. The table shows that our continuous optimization-based method consistently outperforms grid search across benchmarks, independent of cascade length. The only benchmark mark where we report a tie is MMLU; on all other benchmarks, the reductions in AUC are statistically significant at the $p<0.05$ level. On XSum and GSM8K, our method performs especially well, lowering the AUC by 1.5% and 2.5% compared to grid search.
@@
Moving on to a comparison of computational complexity, Figure 6 shows that the runtime of our continuous optimization-based method for finding optimal confidence thresholds scales much more favorably, both in the length of the cascade $k$ as well as the desired resolution $h$ of the cost-error curve. Here, the resolution $h$ refers to the density at which we sample the optimal cost-error curve along the cost-axis. For grid search, $h$ is simply the reciprocal of the number of grid points in each dimension. For our method, $h$ is the reciprocal of the number of times we solve the optimization problem (11). In other words, $h=1/|\Lambda|$, where $\Lambda$ is the set of cost sensitivities we consider in (11).
The favorable scaling behavior of our continuous-optimization based method for selecting confidence thresholds has important implications. Most of all, low-order polynomial scaling in the cascade length $k$ makes it possible to tune confidence thresholds for arbitrarily long cascades. With the proliferation of different LLMs available via third-party APIs, we consider the ability to construct long cascades to be a significant advantage. This prospect becomes especially interesting for cascade models that allow skipping over intermediate models within the cascade, so that inference costs do not sum along the length of the cascade.
@@ 5 Conclusion
Importantly, our probabilistic model yields expressions for a cascade’s probability of correctness and expected cost that are differentiable with respect to the confidence thresholds, making continuous optimization possible. Compared to selecting confidence thresholds using grid search, our continuous-optimization based approach presents significantly enhanced computational scaling, turning the dependence on cascade length $k$ from intractable into low-order polynomial. In addition, the optimal thresholds found using our method increasingly outperform those selected via grid search as cascade length grows: for cascades consisting of $k\geq 3$ models, we have observed a 1.9% average decrease in the area under the cost-error curve on the test set; for $k\geq 4$, the gap widens to 2.2%; and for $k=5$, it stands at 2.6%.
Overall, our results point to a larger vision for the future of deploying LLMs. Using probabilistic models, we will be able to adaptively select the most suitable model to answer each query, improving both reliability and performance. Additionally, probabilistic modeling will make it possible to anticipate the performance of an LLM system under different conditions, making it possible to seamlessly adapt the system as conditions shift. We are excited to further pursue this line of research in subsequent work.
@@ 5.6 TruthfulQA
$\boldsymbol{\Delta}$ Intelligence Provider Smaller Model Larger Model Price Ratio

### v4

@@ 4.4 Rational Tuning of Confidence Thresholds
Table 7: Area under the error-cost curve (AUC) on the test set, showing that our Rational Tuning (“RT”) framework for selecting confidence thresholds consistently outperforms both a Bayesian optimization baseline (“BO”) and high-resolution grid search (“GS”). The mean percentage changes (%$\%\Delta$) are statistically significant at the $p<0.05$ on almost all benchmarks, as measured by Wilcoxon rank-sum tests paired by cascade (highlighted in bold).
In this section, we examine the performance and runtime scaling of our continuous optimization-based algorithm (11) for selecting optimal confidence thresholds. We consider all 26 possible cascades of length $k\geq 2$ composed of Meta’s Llama models (1B, 3B, 8B, 70B, and 405B). We evaluate against Bayesian optimization and high-resolution grid search baselines on six benchmarks (MMLU, MedMCQA, XSum, TriviaQA, GSM8K, TruthfulQA) spanning general-purpose knowledge and reasoning, domain-specific QA, text summarization, open-ended QA, mathematical reasoning, and the ability to avoid hallucinations on adversarial questions.
[…]
In addition, we measure how the runtime for finding optimal confidence thresholds scales with the length of the cascade and the desired resolution of the error-cost curve on the $x$ axis, i.e., how densely we sample the optimal thresholds. We have not overly optimized our code and mainly aim to contrast asymptotic scaling behavior.
Bayesian optimization baseline: this baseline runs Bayesian optimization with a Gaussian process surrogate function, via the HEBO package (Cowen-Rivers et al.,, 2022, Shahriari et al.,, 2016). The Bayesian optimization minimizes (11), in an analogous manner to our Markov-copula (“Rational Tuning”) approach. We run HEBO for as many iterations as needed until the change in loss between successive iterations is below a numerical tolerance ($\epsilon=10^{-5}$). In practice, we found that the final change in loss is typically 0.0. Following the practical guidance of HEBO’s authors555github.com/huawei-noah/HEBO/tree/master/HEBO. Accessed April 6, 2025., we use four parallel suggestions during each iteration. We adaptively interpolate the optimal thresholds computed by HEBO in the same way we do for Rational Tuning (see Equation (13)).
High-resolution grid search baseline: this baseline selects optimal confidence thresholds by searching over adaptive grids computed from the model-specific quantiles of calibrated confidence. Specifically, in each dimension the grid ranges from $\phi_{\text{min}}$ to $\phi_{\text{max}}$ in increments of $2.5\%$ probability mass. This results in considering $40^{k-1}$ candidate threshold combinations for cascades with $k$ models, ranging from 40 candidates for a two-model cascade to $40^{4}$ = 2,560,000 candidates for a five-model cascade. After scoring all candidate threshold combinations, we use the Skyline operator (implemented in the Python package paretoset [^6] Open-source implementation available at github.com/tommyod/paretoset. Accessed January 13, 2025.) to filter the candidate threshold vectors down to the Pareto-optimal set (Börzsönyi et al.,, 2001). A candidate threshold vector $\bm{\theta}=(\phi_{1},...,\phi_{k-1})$ is Pareto-optimal if its performance metrics $(\mathbb{P}_{\bm{\theta}}(\text{Correct}),\mathbb{E}_{\bm{\theta}}[\text{Cost}])$ are not dominated by any other candidate threshold vector $\bm{\theta^{\prime}}$ in the sense that $\mathbb{P}_{\bm{\theta^{\prime}}}(\text{Correct})>\mathbb{P}_{\bm{\theta}}(\text{Correct})$ and $\mathbb{E}_{\bm{\theta^{\prime}}}[\text{Cost}]<\mathbb{E}_{\bm{\theta}}[\text{Cost}]$.
Figure 6 shows that our Rational Tuning framework for selecting confidence thresholds results in lower AUC on the test set compared to the baselines. Each point on the plot shows the average percentage reduction in AUC for all cascades of a given length $k$, averaged across all benchmarks. As cascade length $k$ grows, our method outperforms the baselines by a larger margin. For example, the mean reduction in AUC compared to Bayesian optimization is 4.3% for $k\geq 3$; 5.8% for $k\geq 3$; and 7.2% for $k=5$. The corresponding performance gains relative to high-resolution grid search are 2.0%, 2.2%, and 2.7%. We computed statistical significance of these percentage differences using a Wilcoxon rank-sum test paired by cascade.
We hypothesize that the performance gains of Rational Tuning relative to Bayesian optimization stem from the fact that our framework applies a (mostly correct) inductive assumption about the correlation structure of LLM cascades, whereas Bayesian optimization is a general-purpose algorithm for black-box optimization. Section 4.4.1 corroborates this hypothesis by presenting more significant performance gains in the low-sample limit with $n\leq 30$ training examples.
By contrast, it is not surprising that grid search performs worse as $k$ increases, since searching for effective threshold combinations by trial and error suffers from the curse of dimensionality.
Table 7 presents the results broken down by benchmark rather than cascade length. The table shows that Rational Tuning consistently outperforms Bayesian optimization and grid search across benchmarks, independent of cascade length. The only benchmark mark where we report a tie is MMLU. On almost all benchmarks, the reductions in AUC are statistically significant at the $p<0.05$ level.
4.4.1 Performance in the Low-Sample Limit
Our Rational Tuning methodology relies on a small labeled training data set consisting of LLM confidence scores and corresponding binary correctness labels. Since labeled data is scarce in many applications, we supplement our main experiment ($n\approx 300$ training examples) with a study of the low-sample limit. Here, we re-run our experiment for $n\leq 30$ training examples. For each benchmark, we ensure a balanced subsample with both correct and incorrect answers from each model by sampling training examples in pairs (one correct, one incorrect for each model) for a fixed number of iterations, with a target of $n=30$ examples. Since collisions may occur, the final number of sampled training examples lies between 20 and 30, depending on the benchmark.
Figure 7 displays the results, revealing that Rational Tuning significantly outperforms the Bayesian optimization and grid search baselines as cascade length grows. On cascades with $k\geq 3$ models, the average performance gain is 10.2% relative to Bayesian optimization, and 5.6% relative to high-resolution grid search.
Table 8 breaks down these results by benchmark. We see that Rational Tuning outperforms the baselines on each benchmark except for TruthfulQA, where high-resolution grid search performs best. On the other benchmarks (MMLU, MedMCQA, TriviaQA, XSum, and GSM8K), the performance gains of our method are statistically significant at the $p<0.05$ level, according to Wilcoxon rank-sum tests paired by cascade.
Our interpretation of these results is that our Rational Tuning framework benefits from making inductive assumptions about the interactions between the error rates of different LLMs. Crucially, fitting these inductive assumptions to empirical data requires only few observations: since each copula model depends on a single scalar correlation parameter $\theta\in\mathbb{R}$, our method requires only $k-1$ parameters to model the interactions between the error rates of $k$ different LLMs.
(a) Relative to Bayesian optimization
(b) Relative to grid search
Figure 7: Reduction in the area under the error-cost curve (AUC) as cascade length grows, in the low-sample limit ($n\leq 30$ training examples), when using our Rational Tuning framework. In (a), we compare against a Bayesian optimization baseline, while in (b) we compare against high-resolution grid search. Our method increasingly outperforms the baselines as cascade length grows. Error bars show ±$\pm 1\sigma$ of the mean percentage change, and filled markers indicate statistical significance.
Benchmark AUC $\bm{\downarrow}$ RT (Ours) vs GS RT (Ours) vs BO
RT (Ours) GS BO %$\%\Delta_{\mathrm{GS}}\downarrow$ $p_{\mathrm{GS}}$$\%\Delta_{\mathrm{BO}}\downarrow$ $p_{\mathrm{BO}}$
MMLU $\bm{0.293}$ 0.308 0.316 -5.51 $\bm{8.0\times 10^{-6}}$7.74 $\bm{1.0\times 10^{-6}}$
MedMCQA $\bm{0.399}$ 0.416 0.419 -4.18 $\bm{1.1\times 10^{-5}}$4.72 $\bm{5.0\times 10^{-6}}$
TriviaQA $\bm{0.197}$ 0.200 0.203 -2.11 $\bm{1.5\times 10^{-2}}$3.26 $\bm{5.1\times 10^{-3}}$
XSum $\bm{0.422}$ 0.431 0.442 -2.42 $\bm{2.6\times 10^{-2}}$5.06 $\bm{4.2\times 10^{-4}}$
GSM8K $\bm{0.187}$ 0.195 0.197 -3.41 $\bm{2.5\times 10^{-2}}$4.21 $\bm{2.1\times 10^{-4}}$
TruthfulQA 0.449 $\bm{0.442}$ 0.452 1.60 $1.0\times 10^{0}$0.73 $1.2\times 10^{-1}$
Average $\bm{0.325}$ 0.332 0.338 -2.67 -4.29
Table 8: Area under the error-cost curve (AUC) in the low-sample limit ($n\leq 30$ training examples), showing that our Rational Tuning (“RT”) framework for selecting confidence thresholds consistently outperforms both a Bayesian optimization baseline (“BO”) and high-resolution grid search (“GS”). The mean percentage changes (%$\%\Delta$) are statistically significant at the $p<0.05$ on almost all benchmarks, as measured by Wilcoxon rank-sum tests paired by cascade (highlighted in bold).
4.4.2 Sensitivity to Statistical Assumptions
(a) Copula $\sqrt{n}\text{CvM}$
(b) Marginal $\sqrt{\text{CvM}}$
Figure 8: Sensitivity of Rational Tuning’s performance gains to the Cramér-von Mises (CvM) test statistics (lower is better). Overall, performance appears to be more sensitive to mis-specification of the copula model.
Our implementation of Rational Tuning models uses mixtures of beta distributions to model the marginal distribution of confidence scores, and Gumbel copulas to model pairwise correlations between LLMs. In Section 4.3, we quantify the deviation between these modeling assumptions and the true empirical distributions via Cramér-von Mises (CvM) statistics.
Figure 8 shows the sensitivity of Rational Tuning’s performance gains (relative to the high-resolution grid search baseline) to the mean CvM statistics for each cascade. For example, if $M_{1}\rightarrow...\rightarrow M_{k}$ has marginal $\sqrt{\text{CvM}}$ scores $\sigma_{1},...,\sigma_{k}$ and copula $\sqrt{n}\text{CvM}$ scores $\sigma_{1,2},\sigma_{2,3},...,\sigma_{k-1,k}$, then the mean marginal and copula CvM scores are $\overline{\sigma}_{\text{marginal}}=\frac{1}{k}\sum_{i=1}^{k}\sigma_{i}$, and $\overline{\sigma}_{\text{copula}}=\frac{1}{k}\sum_{i=2}^{k}\sigma_{i-1,i}$, respectively.
Figures 8a and 8b suggest that lower CvM divergences improve the relative performance of Rational Tuning. This effect is more pronounced for the copula statistics rather than the marginal statistics, highlighting the importance of correctly modeling the correlations between LLMs. In both plots, the light blue data points with little performance gain despite excellent CvM values are heavily enriched for two-model cascades. For cascades with $k=2$ models, Rational Tuning generally performs on par with grid search or Bayesian optimization.
The plots also suggest some robustness to deviations from the statistical assumptions. We are eager to explore Rational Tuning’s robustness to model mis-specification in greater detail in subsequent work.
4.4.3 Computational Scaling
@@
(b) The runtime of our method always grows linearly in the desired resolution $h$ at which the error-cost curve is sampled along the cost axis, independent of the cascade length $k$. However, the runtime of grid search scales as $h^{k-1}$ in $h$ and $k$ (log-log plot).
@@
Figure 9: Shows runtime scaling for computing the full error-cost curve, comparing our continuous-optimization based algorithm (“continuous”, blue) to grid search (“grid search”, gray). Our method scales much more favorably as the cascade length grows, and as the error-cost curve is sampled more densely along the cost axis. The shading shows ±$\pm 1\sigma$ of the observed data points.
@@
Moving on to a comparison of computational complexity, Figure 9 shows that the runtime of our Rational Tuning framework for finding optimal confidence thresholds scales much more favorably compared to grid search, both in the length of the cascade $k$ as well as the desired resolution $h$ of the error-cost curve. Here, the resolution $h$ refers to the density at which we sample the optimal error-cost curve along the cost-axis. For grid search, $h$ is simply the reciprocal of the number of grid points in each dimension. For our method, $h$ is the reciprocal of the number of times we solve the optimization problem (11). In other words, $h=1/|\Lambda|$, where $\Lambda$ is the set of cost sensitivities we consider in (11).
We omit Bayesian optimization from Figure 9, since we observed longer runtimes (10-1000x longer) that exhibit less clear scaling with the length $k$ of the cascade. Specifically, the average number of iterations until convergence increases from 2.9 for $k=2$ to 5.3 for $k=4$, then drops back to 4.0 for $k=5$. Across the data, the minimum and maximum number of iterations required until convergence are 2 and 14.
4.4.4 Practical Guidelines
Our experiments suggest that Rational Tuning is the preferred methodology for tuning confidence thresholds for longer cascades ($k\geq 3$) with multiple deferral thresholds. However, for two-model cascades ($k=2$) with a single deferral threshold, the Markov assumption is void (as there are only two models). In this setting, the nonparametric nature of one-dimensional grid search should give the most reliable results.
When applying Rational Tuning for cascades with $k\geq 3$ models, we recommend visually inspecting the match between the assumed probabilistic model and the empirical data. Specifically, we recommend the following visual diagnostics:
  • Marginal distributions: compare histograms of the fitted and empirical distributions (as in Figure 4) and verify that the overall fit is adequate.
  • Pairwise correlations: construct copula plots as in Figures 2 and 3. Compare these plots to random samples from a Gumbel copula with correlation parameter $\hat{\theta}=\frac{1}{1-\hat{\tau}}$, where $\hat{\tau}$ is the empirical rank correlation.
  • Markov assumption: construct rank correlation plots as in Figure 1. Verify that rank correlations are strong near the diagonal.
In addition, it is important to assess the expected calibration error (ECE) of the confidence scores. We recommend computing the ECE using quantile binning with 10 or 20 bins; ideally, the ECE should not exceed $10\%$.
@@ 5 Conclusion
Importantly, our probabilistic model yields analytical expressions for a cascade’s error rate and expected inference cost. These expressions are differentiable with respect to the cascade’s confidence thresholds, making continuous optimization possible. Compared to selecting confidence thresholds using Bayesian optimization and high-resolution grid search, our Rational Tuning framework yields more favorable error-cost trade-offs as cascade length grows, outperforming the baselines by up to 7.2% when using $n\approx 300$ labeled training examples. In the low-sample limit ($n\leq 30$ training examples), the performance gains reach up to 16.5%, suggesting that our framework’s inductive assumptions about the interactions between the error rates of different LLMs improve sample efficiency.
Building on these promising results, an interesting direction would be to apply our probabilistic modeling framework to LLM routing, in which a central routing model sends a query to the most suitable LLM in a single step, avoiding cumulative cost increases as the query propagates down a cascade. Since cumulative cost increases are especially severe for longer cascades (at which our methodology excels), the routing setting may more effectively leverage Rational Tuning’s capacity for modeling dependencies between arbitrarily many distinct LLMs. For instance, suppose the routing decision depends on noisy estimates $\hat{\phi}_{1},...,\hat{\phi}_{n}$ of the LLMs’ true calibrated confidences. In this case, balancing the noisy observations $\hat{\phi}_{i}$ against their probabilistic expectations $\mathbb{E}[\hat{\phi}_{i}|\hat{\phi}_{1},...,\hat{\phi}_{i-1},\hat{\phi}_{i+1},...\hat{\phi}_{n}]$ may lead to more effective routing decisions.
Ultimately, our results point to a larger vision for the future of deploying LLMs. Using probabilistic models, we will be able to adaptively select the most suitable model to answer each query, improving both reliability and performance. Additionally, probabilistic modeling will enable us to anticipate the performance of a system of LLMs under different conditions, making it possible to seamlessly adapt the system as conditions shift. We are excited to further pursue this line of research in subsequent work.
@@ 5.6 TruthfulQA
$\bm{\Delta}$ Intelligence Provider Smaller Model Larger Model Price Ratio
@@
6 Appendix F: Recomputing Rank Correlations on Correct and Incorrect Answers
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 11: MMLU: Kendall’s $\tau$ rank correlations of Llama3 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 12: MMLU: Kendall’s $\tau$ rank correlations of Llama3, GPT-4o, and Qwen2.5 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 13: MedMCQA: Kendall’s $\tau$ rank correlations of Llama3 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 14: MedMCQA: Kendall’s $\tau$ rank correlations of Llama3, GPT-4o, and Qwen2.5 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 15: TriviaQA: Kendall’s $\tau$ rank correlations of Llama3 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 16: TriviaQA: Kendall’s $\tau$ rank correlations of Llama3, GPT-4o, and Qwen2.5 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 17: XSum: Kendall’s $\tau$ rank correlations of Llama3 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 18: XSum: Kendall’s $\tau$ rank correlations of Llama3, GPT-4o, and Qwen2.5 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 19: GSM8K: Kendall’s $\tau$ rank correlations of Llama3 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 20: GSM8K: Kendall’s $\tau$ rank correlations of Llama3, GPT-4o, and Qwen2.5 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 21: TruthfulQA: Kendall’s $\tau$ rank correlations of Llama3 models ordered by size.
(a) Incorrect Only
(b) Correct Only
(c) All
Figure 22: TruthfulQA: Kendall’s $\tau$ rank correlations of Llama3, GPT-4o, and Qwen2.5 models ordered by size.
In this section, we verify the rank correlations between the confidence scores of different LLMs by recomputing them conditioned on both models answering correctly or incorrectly.
Figures 11-22 extend Figure 1 by separately re-computing the rank correlation patterns on correctly and incorrectly answered queries. In addition, Table 10 below shows the average rank correlations computed separately on the correct, incorrect, and all answers, for each benchmark. We compute $\tau_{\text{inc}}$, $\tau_{\text{corr}}$, $\tau_{\text{all}}$ for each pair of models, as well as the rank correlations between these measurements across model pairs: $\tau_{\text{inc, corr}}$, $\tau_{\text{inc, all}}$, $\tau_{\text{corr, all}}$.
Note that since error rates are low for some models and benchmarks (see Table 1), conditioning on incorrectly answered queries leaves only few observations for some model pairs. In Figures 11-22, we print “?” for rank correlations with sample size less than 50; we use the $n=50$ cut-off since it reduces the standard error for Kendall’s $\tau$ to around $\sigma_{\tau}\leq 0.1$, based on a normal approximation.
Table 10: Rank correlations of calibrated confidences between Llama3 1B, 3B, 8B, 70B, and 405B models, computed separately for incorrectly answered queries (“inc”), correctly answered queries (“corr”), and all queries (“all”): $\overline{\tau}$ is the average rank correlation across the 10 model pairs; $\tau_{a,b}$ is the rank correlation between data subsets (inc, corr, all) across model pairs; and $p_{a,b}$ is the $p$ value for $\tau_{a,b}$.
Benchmark $\bm{\overline{\tau}_{\text{inc}}}$ $\bm{\overline{\tau}_{\text{corr}}}$ $\bm{\overline{\tau}_{\text{all}}}$ $\bm{\tau_{\text{inc,corr}}}$ $\bm{p_{\text{inc,corr}}}$ $\bm{\tau_{\text{inc,all}}}$ $\bm{p_{\text{inc,all}}}$ $\bm{\tau_{\text{corr,all}}}$ $\bm{p_{\text{corr,all}}}$
MMLU 0.199 0.396 0.385 0.505 $<0.0001$ 0.717 $<0.0001$ 0.686 $<0.0001$
MedMCQA 0.174 0.295 0.271 0.324 0.0055 0.486 $<0.0001$ 0.730 $<0.0001$
TriviaQA 0.110 0.274 0.298 0.371 0.0014 0.413 0.0004 0.819 $<0.0001$
XSum 0.138 0.049 0.065 0.180 0.1284 0.187 0.1148 0.721 $<0.0001$
GSM8K 0.199 0.169 0.201 0.425 0.0003 0.467 $<0.0001$ 0.921 $<0.0001$
TruthfulQA 0.222 0.196 0.178 0.667 $<0.0001$ 0.860 $<0.0001$ 0.756 $<0.0001$